### v1

@@ Abstract
Many automated decision systems (ADS) are designed to solve prediction problems— where the goal is to learn patterns from a sample of the population and apply them to individuals from the same population. In reality, these prediction systems operationalize holistic policy interventions in deployment. Once deployed, ADS can shape impacted population outcomes through an effective policy change in how decision-makers operate, while also being defined by past and present interactions between stakeholders and the limitations of existing organizational, as well as societal, infrastructure and context. In this work, we consider the ways in which we must shift from a prediction-focused paradigm to an interventionist paradigm when considering the impact of ADS within social systems. We argue this requires a new default problem setup for ADS beyond prediction, to instead consider predictions as decision support, final decisions, and outcomes. We highlight how this perspective unifies modern statistical frameworks and other tools to study the design, implementation, and evaluation of ADS systems, and point to the research directions necessary to operationalize this paradigm shift. Using these tools, we characterize the limitations of focusing on isolated prediction tasks, and lay the foundation for a more intervention-oriented approach to developing and deploying ADS.
@@ 1   Introduction
We focus on prediction-based interventions in social systems, specifically automated decision systems used for institutional decisions that impact human well-being and opportunity. These systems predict (latent or future) outcomes to support decision-making. Certainly, there are many different sociotechnical arenas where predictions impact individuals, beyond our focus on ADS in consequential social systems: such as online platforms, management of complex systems, or social media. In comparison to these other venues, for example, we focus on the introduction of ADS into existing social processes: decision-makers interact with the algorithm’s recommendations and have agency to act, while decision subjects can only respond, appeal, or contest them. Typically, there are far fewer decision-makers than decision subjects, as seen in examples like teachers and students or doctors and patients. as such, the users of ADS products we consider for this paper are institutional decision makers rather than individual consumers. In our context of study, the algorithm is operationalizing a policy change within an institutional context – at times, mediated or defined by a broader societal or government policy, but not necessarily so.
@@ 2   Background
We argue that typically, introducing a new predictive model in an intake or targeting process is a programmatic change that affects decision outputs, and therefore outcomes. Outcomes include both short-term, measurable outcomes as well as medium- and longer-term outcomes and overall policy goals (such as improving health outcomes broadly).
In turn, formulating, developing, and maintaining a predictive model requires marshaling many of an organization’s resources, which we broadly conceptualize as the total cost of ownership of ADS. Organizations in consequential domains are constantly tasked with doing more with less, and they need to use these limited resources to develop, procure or pay for ADS, potentially instead of other programmatic changes. Is it worth it, in terms of improving societal outcomes? It is urgent to establish a careful science of evaluating ADS’ impacts to actively steer their development, rather than leaving it up to hype or inertia.
@@
Beyond the institutional context of a single organization and its activities, predictive models are deployed in broader policy contexts, which shape their impacts in important ways. Systems incorporating predictive models may optimize a status quo policy arrangement, or lend legitimacy to and operationalize austerity policies and cost-cutting measures that maintain resource scarcity. Their introduction may also often coincide with other policy changes and aligned reform movements, making it difficult to isolate the specific effects of the predictive tools themselves.
For example, for risk assessment instruments (RAIs) in criminal justice, it is difficult to disentangle success stories of RAIs achieving decarceral aims, such as reducing jail populations, from broader movements of prosecutorial reform. Conversely, attempts to enact reforms by computational changes to ADS are brittle: if front-line decision-makers disagree with the goals, they might simply ignore ADS information.
Example 1 (Risk assessment for pretrial detention).
Pretrial risk assessments aim to predict a defendant’s failure to appear for the court date or committing a new crime. These risk assessment tools are used to inform pretrial detention decisions from judges, as they decide among different options: releasing someone pretrial without conditions, release with conditions (like supervision or other mandated programs), setting a money bail amount, or pretrial detention. A defendant who is detained pretrial remains in jail before their trial. Since defendants are presumed innocent before proven guilty, detention should be reserved for those at greatest risk of failing to appear or committing another crime.
Several policy proposals exist in order to minimize the reliance on cash bail in various U.S. states. Most notably, New Jersey in 2017 voted to replace the discretion of judges with an algorithmic assessment of the recidivism risk of defendants. Since its widespread deployment, pretrial detention incarceration has decreased by 67% (Anderson et al., 2019). On the other hand, California voted against a similar proposal, Proposition 25, with advocacy groups citing concerns around the harms possibly posed by the inaccuracy and bias reported in algorithmic risk assessments (Green, 2018).
@@
Therefore, current decisions made in these organizations reflect a combination of institutional policy, categorical prioritization, and human discretion. In Figure 1, this is represented via the database of prior decisions and outcomes ($D$,$Y$). In the next sections, we focus on formal problem formulations to catalog predictions vs. interventions.
@@
Institutional context and their ongoing practices introduce idiosyncratic variability in the requirements and goals of shared predictive tasks like “predict heart disease” or “predict recidivism”. A general task like “predict heart disease” becomes “predict heart disease when doctors and hospitals have protocol A,B,C for tests and medications, and the patient case mix is D.”
@@
Institutional context dictates what data is collected in the first place. For example, in Example 2 (cost prediction), a machine learning researcher might consider cost prediction rather than outcome prediction, as in (Obermeyer et al., 2019), to be an error of naivete rather than a natural consequence of how institutions generate data. A lot of data collection in healthcare is to support insurance billing, and was not collected in the first place to support downstream predictive innovation or future machine-learning papers. Most data from institutions is collected for administrative purposes for their ongoing activities.
@@ 3   Model Design
Although predictive risk models have a simple formulation—predict $Y$ from $X$—they in fact are used in a wider variety of ways to impact and change future real-world outcomes.
ADS directs additional resources, scrutiny, services, protection/prevention, or sanction. These decisions often use individual-level information to make targeted decisions and need not be data-driven. These distinct use cases of ADS often imply different appropriate technical formulations beyond pure prediction. In criminal justice (Example 1), pretrial risk assessment tools and their predictions of failure to appear are used to inform decisions about pretrial detention or release. But the same predictions are also used to recommend a different intervention of supervised release, which can comprise beneficial services or additional surveillance. In housing (Example 3), the VI-SPDAT assesses vulnerability and housing need; it is associated with returns to homelessness (but evidence is mixed). The task of predicting returns to homelessness is however different from assessing potential differential impacts of differing levels of housing intensity.
@@
  • For an individual with covariates $X=x$, and given a threshold $t$, a binary classifier can be defined as $\hat{Y}=\mathbb{I}[\hat{R}(x)>t]$, $\hat{Y}\in\{0,1\}$, with natural extensions to multi-class or other structured outputs via analogous thresholding rules.
@@
While we use $D$ to represent historical decisions, we consider alternative decision-making procedures, summarized in the decision policy $\pi(X):\mathcal{X}\mapsto\mathcal{D}$. Predictive risk modeling is a common tool that is often used to triage scarce resources. When used as such, often the decisions are thresholded based on predictive risk,
@@
In practice, we find that purely-predictive formulations that target on $\mathbb{E}[Y\mid X]$ are more common, i.e. learning a predictive model based on all prior data without any explicit modeling of prior decisions and interventions. However, if prior data does includes other decisions or interventions, this estimand hides the resulting heterogeneity, and is vulnerable to distribution shifts in interventions (including those across jurisdictions when an ADS developed in one is used in another). For example, a prior pilot study may not continue or might in fact scale up. Learning $\mathbb{E}[Y\mid X]$ from the pooled historical data would be brittle to such shifts, while the counterfactual risk models discussed above would remain robust. Koepke and Robinson (2018) argue that ignoring recent reforms or other changes in the social world could lead to “zombie predictions” that overestimate risk. These interventions in the historical data include pretrial services that reduce rearrest risk or bail reforms that reduce pretrial detention and iatrogenic rearrests.
@@
Intervention-aware targeting offers many different options beyond predictive risk models of $\mathbb{E}[Y\mid X]$ alone. Recent discussions of predictions vs. interventions include Barabas et al. (2018), which argues that the use of predictive risk models is a poor fit for improving social conditions, drawing on the case of risk assessments in criminal justice.666 Barabas et al. (2018) argues that predictive models that focus on merely predicting future outcomes neither manage long-term goals of reducing reoffense (especially when interactions with the criminal justice system are iatrogenic, increasing criminal activity), nor reduce criminal risk, nor improve understanding of root causes of criminal behavior that could be used to design interventions to reduce risk. This critique, among others, points out the urgency of going beyond prediction to consider alternative avenues for improving social outcomes. Prediction by itself doesn’t directly achieve long-term goals or inform interventions to reduce risk, as argued in Liu et al. (2023a). One can also reconceptualize predictive risk models in terms of the subjects of prediction. [^7] Meyer et al. (2022) highlight how current risk prediction focuses narrowly on harms individuals may pose to systems (e.g., future crime); by “flipping the script,” they show how prediction could just as well be used to predict system harms to individuals, such as lengthy sentences. So much more is possible beyond current formulations.
@@
Understanding the institutional context is key to inform prioritization for supervised release. Pretrial detention can preclude re-offense by its nature, albeit at a large cost to civil liberties. On the other hand, supervised release is a much “softer” decision and its effects on re-offense are unclear. How should judges determine who should receive supervised release? There is some broad guidance that “release conditions—if any are considered necessary—should be the least restrictive that reasonably assure court appearance and community safety,” (APPR, 2024) emphasizing $Y(D)$, the final outcomes of whatever release conditions are provided. Other guidance does indicate some preference for reducing risk of the riskiest individuals, suggesting that “[limited community] resources should go to people with the highest need, meaning people assessed as the least likely to succeed.” Heterogeneous treatment effects can introduce tradeoffs between these two competing objectives. Increased surveillance can increase frequency of technical violations of the supervision conditions and therefore worsen outcomes. Supportive services might work for some on the margin, but may not be enough to overcome other structural barriers. The limited evidence base is mixed, and variation in its program design makes generalization from prior studies even more difficult.
@@
recommended $\text{recommended supervision}=\mathbb{I}\left[\mathbb{E}[Y^{FTA}(0)+Y^{NCA}(0)\mid X]>t\right],$
@@
We discuss details of the study in more detail in Appendix B. Even though the risk prediction model of Black et al. (2003) improved predictions compared to prior models in other states, in part by using additional covariates and using a continuous rather than binarized outcomes, they find that predicted duration is in general unrelated to heterogeneous treatment effects. In this setting, equity vs. efficiency objectives might favor different problem formulations. Assuming that long-term spells on unemployment reflect extended precarity and reflect the worst-off, proritizing long-term spells in the absence of services might reflect attention to providing services to the worst-off. But from the point of view of the UI system, targeting services to those most likely to benefit increases efficient use of resources, and could expand capacity. Improving outcomes would require changing the allocation mechanism and deliberating these different trade-offs, beyond just improving prediction.
Other recent studies assess targeting based on baseline risk or need vs. welfare-efficient ITRs in different domains. Haushofer et al. (2022) compare typical eligibility for cash transfers by proxy-means testing (predicting household deprivation based on simple observable characteristics, such as having a thatched roof) vs thresholding on HTEs. Athey et al. (2025) compare targeting nudges for completing financial aid applications based on predicted baseline outcomes vs. HTEs, and also develop hybrid approaches. Sirota et al. (2025) compare predictive risk targeting vs. individual treatment regimes in healthcare, for cardiovascular heatlh.
Finally, it is important to recognize that there can be other factors that justify targeting on $Y(0)$ rather than $Y(1)-Y(0)$. For certain problem domains, there can be strong moral and ethical reasons to prioritize benefit-based interventions based on baseline risk rather than average social welfare alone. Prioritization based on baseline risk can align with intuitive or politically palatable conceptions of who deserves help. [^11] Gowan (2010) studies homelessness and policy developments in San Francisco; the medicalization of homelessness received greater institutional support and influenced the development of vulnerability assessments that focus on mortality risk. In turn, these vulnerability assessments individualize homelessness and emphasize individual-level health factors rather than broader structural and systemic factors contributing to homelessness. Although we haven’t discussed estimation in detail, it is much harder to estimate $Y(1)-Y(0)$ than the baseline risk, and impossible to evaluate in prior data (as we discuss in Section 4). We discuss how the statistical limits of prediction might inform problem formulation in Section 3
@@
Formally, the first case is where we are able to specify a loss function $\ell(D,Y)$ which depends on an observable quantity $Y$ which is not itself a function of $D$, e.g., the decision to carry an umbrella does not itself impact whether it rains. The second case is where the appropriate loss function is parameterized by quantities that are not fully observable due to decision-dependence. For patients that we do not treat, we cannot observe their treatment effects, and vice versa for those we do treat.
$\ell(D,Y)$
@@
What are examples where prediction is enough? The “umbrella” dynamic can arise when targeting programs that might act as social insurance, therefore kicking in with high efficacy conditional on random events (to be predicted). An example is intensive homelessness prevention.
Example 6 (Housing - Preventing Homelessness (California Policy Lab) ).
The California Policy Lab and the University of Chicago Poverty Lab have used multi-system service utilization data in Los Angeles county in order to predict homelessness, in particular to target outreach with cash benefits to try to predict and prevent the onset of homelessness (Wachter et al., 2019).
Pure prediction if the outreach (a sizeable subsidy to spend on whatever needs are most pressing) is extremely effective in preventing homelessness — the magnitude of the effect outweighs heterogeneity concerns. [^12] Liu et al. (2023a) formalized this notion of a sufficient action that improves outcomes for all decision subjects, regardless of their particular circumstances. Using a causal-graphical formalism, they show that the existence of such an action is both necessary and sufficient for pure prediction to be adequate for targeting interventions.
@@
We now turn to learning decision policies for the “pure prediction” setting (fig. 5(a)). Formally, this is where $\ell$ depends only on an outcome $Y$ that is always observable for historical training samples. That is, we observe samples $(X_{1},Y_{1})...(X_{n},Y_{n})\sim P$. The basic idea is a “plug-in” strategy to minimize expected loss. Proceeding conditionally on a particular covariate value $x\in\mathcal{X}$
[…]
The plug-in strategy first estimates the unknown conditional probability density ${p}(y\mid x)=P(Y=y\mid X=x)$, and then plugs-in the estimated ${p}(y\mid x)$ to evaluate the objective and optimize $\pi$.
This is optimal when plugging in the true conditional probability. But this is typically unattainable in real-world prediction, motivating alternative decision-aware methodologies. How do we make the best decision possible from an imperfect predictive model?
Decision rules.
One strategy collapses the two stages of the plugin strategy: instead of first predicting $Y$ from $X$ and then choosing $D$ based on $E[Y\mid X]$, we could simply learn a mapping from $X$ to $D$. One common way of doing so is to optimize the decision rule over a parameterized class of models $\Pi$.
$\pi^{\star}=\arg\min_{\pi\in\Pi}\mathbb{E}[\ell(\pi(X),Y)]$
The above can be estimated on data. Directly learning the decision rule reconciles trade-offs in favor of the loss function. There is an extensive literature on such approaches, particularly in the operations research community.
The decision rule formulation is typically motivated by settings where two properties hold: (1) structure of the action space is relatively simple (e.g., a small, discrete set of actions) and (2) the optimal mapping $X\to D$ is believed to not be too much more complex than the optimal mapping $Y\to D$.
Decision-aware learning
Another more advanced formulation seeks to learn predictive models that explicitly achieve the best decision-loss. Decision-aware methods seek a decision-aware predictor function $f_{\mathrm{DA}}(y\mid x)$ which achieves the best expected loss, under the decisions optimal for $f_{\mathrm{DA}}$, denoted as $\pi^{*}_{f}(X)$.
[…]
See Appendix B for more details. We mention this more advanced approach to highlight the importance of articulating how predictors affect downstream decisions. An example of this type of non-trivial decision-loss structure arises when cost minimization relies on global system state, such as shortest-paths, routing, and/or supply chain problems.
@@
In contrast, decisions impact outcomes in unknown ways (fig. 5(b)), targeting decisions requires estimating HTEs, as we have discussed before. For example, if our goal is to optimize post-decision outcomes:
$\min_{\pi\in\Pi}\mathbb{E}[Y(\pi(X))]$
then the implicit loss function $\ell(D,(Y(0),Y(1)))=Y(D)$ depends on both $Y(1)$ and $Y(0)$, which are never simultaneously observable.
@@
Although we have laid out basic distinctions between predictions and decisions, specifying the objective function and the constraints of the decision problem are also crucial. We refer the interested reader to (Barocas et al., 2023) and references therein for discussions of additional desiderata, such as algorithmic fairness.131313 The algorithmic fairness literature has developed tools for formalizing certain desiderata (such as group or individual fairness), and achieving them, often with constrained optimization. In practice, while auditing for unfairness has surfaced harms and issues with deployed predictive models, sources of unfairness and appropriate mitigations vary widely: perhaps an inappropriate target variable was chosen, the data ecosystem introduces statistical bias, the prediction only informs an institutional decision that is missing from the problem formulation, and so on. A central critique of considerations in the algorithmic fairness literature—that is closely related to the theme in this work—is the mismatch between constraints on model predictions and actual utility or improvements from decisions (see e.g., Liu et al., 2018). Other concerns include “leveling down” objections to imposing fairness constraints that might worsen outcomes for one group while improving them for others (i.e. Pareto-suboptimal transfers of welfare) (Temkin, 2000; Mittelstadt et al., 2023; Hu and Chen, 2020; Maheshwari et al., 2023). Other perspectives on algorithmic fairness emphasize consequentialist or welfarist accounts (Corbett-Davies et al., 2017; Corbett-Davies and Goel, 2018; Hu and Chen, 2020), which are well aligned with our focus on evaluating realized impacts. Accurate and contextual measurement of realized impacts is a crucial input into any of these perspectives.
@@
While we have outlined different technical formulations, how does one translate a real-world problem into a choice for a particular approach? We argue that problem formulation ought to be informed by the ultimate goals of deployment, contested as they may be. Different technical formulations are optimal or appropriate for different objectives, values, and desiderata. Deliberations ought to first take place in the space of the latter – what are we trying to achieve/improve?
We discuss important organizational considerations in problem formulation. Next, we discuss what kind of domain knowledge is required. Finally, even if prediction is sufficient, our prior discussions suggest opportunities to leverage interventional analysis to improve the design of predictive risk models.
@@
Problem formulation for prediction systems is inherently challenging. Effective formulations account for goals, context, decision-structure of prediction-informed processes, and human dynamics of deployment settings.
In data-rich settings, it can be tempting to instead take a ‘data-driven’ opportunistic approach: exploring opportunities to innovate by training candidate predictive risk models on available outcomes in the data, and proceeding further if a model achieves a high enough accuracy. Such an approach, however, is liable to produce algorithms that are ill-suited for the contexts in which they will be used (Coston et al., 2023). Problem formulation should begin before looking at any data, involving key stakeholders in a conversation about the goals, success criteria, and potential risks of an algorithm in the proposed context (Kawakami et al., 2024).
The act of formalizing a problem and building a predictive model is in itself a policy intervention (Abebe et al., 2020). Formalizing a problem can enforce clarity around goals, but introducing a quantitative frame for prediction can distort multiple goals by elevating and operationalizing some but not others (Green and Chen, 2021; Johnson and Zhang, 2022; Liu et al., 2023b).
Efforts to improve the model design and development process would benefit from an understanding of how data science teams currently navigate the model development process. How do data scientists reason about model validity, and which tools, workflows, or resources might best scaffold their processes going forward? Are there any bottlenecks (e.g., training, tools, operational constraints) which limit model development teams’ ability to develop carefully designed systems? While prior studies offer an entry point into these and related questions (Passi and Barocas, 2019; Passi and Jackson, 2018; Guerdan et al., 2025; Kross and Guo, 2021), more research is needed.
Broader community context
Beyond bridging technical perspectives on problem formulation, it’s crucial to recognize inherent asymmetric power dynamics when developing infrastructure for algorithm-informed decisions affecting people (Krafft et al., 2021; Alkhatib and Bernstein, 2019). On the other hand, when it comes to sharing data, data privacy and surveillance concerns are crucial. Actionable Intelligence for Social Policy develop a “Toolkit for Centering Racial Equity Throughout Data Integration”, emphasizing careful consideration of the risks of data integration (disclosure risk, surveillance/sanction risk) against the benefits. This makes it all the more important to evaluate the impacts of ADS interventions on potential benefits. It is crucial to center affected community perspectives. We discuss this and other elements of more holistic design and evaluation of ADS in Section 6.
Formulating evaluations: operationalizing outcomes

### v2

@@ Abstract
Many automated decision systems (ADS) are designed to solve prediction problems— where the goal is to learn patterns from a sample of the population and apply them to individuals from the same population. In reality, these prediction systems operationalize holistic policy interventions in deployment. Once deployed, ADS can shape impacted population outcomes through an effective policy change in how decision-makers operate, while also being defined by past and present interactions between stakeholders and the limitations of existing organizational, as well as societal, infrastructure and context. In this work, we consider the ways in which we must shift from a prediction-focused paradigm to an intervention-oriented paradigm when considering the impact of ADS within social systems. We argue this requires an expansion of the default problem setup for ADS to move beyond prediction to consider the implications of the decision-making context, and downstream outcomes. We highlight how this perspective unifies modern statistical frameworks and other tools to study the design, implementation, and evaluation of ADS systems, and point to the research directions necessary to operationalize this paradigm shift. Using these tools, we characterize the limitations of focusing on isolated prediction tasks, and lay the foundation for a more intervention-oriented approach to developing and deploying ADS.
@@ 1  Introduction
We focus on prediction-based interventions in social systems, specifically automated decision systems used for institutional decisions that impact human well-being and opportunity. These systems predict (latent or future) outcomes to support decision-making. Certainly, there are many different sociotechnical arenas where predictions impact individuals, beyond our focus on ADS in consequential social systems: such as online platforms, management of complex systems, or social media. In comparison to these other venues, for example, we focus on the introduction of ADS into existing social processes: decision-makers interact with the algorithm’s recommendations and have agency to act, while decision subjects can only respond, appeal, or contest them. Typically, there are far fewer decision-makers than decision subjects, as seen in examples like teachers and students or doctors and patients. The users of ADS products are institutional decision makers rather than individual consumers. Often institutions adopt ADS to operationalize a broader policy change, potentially mediated or defined by a broader societal or government policy, but not necessarily so.
@@ 2  Background
What we call the total cost of ownership of ADS refers to the substantial resource investment required to formulate, develop, evaluate and maintain a predictive model and/or ADS. Since organizations in consequential domains are constantly tasked with doing more with less, a key question iswhether the total cost of ownership of ADS is worth it, in terms of improving societal outcomes, compared to alternatives. It is urgent to establish a careful science of evaluating ADS’ impacts so that institutions can actively steer ADS development, rather than being steered by hype or inertia.
@@
Beyond the institutional context of a single organization and its activities, predictive models are deployed in broader policy contexts, which shape their impacts in important ways. ADS may enforce status quo policy by another hand, or lend legitimacy to and operationalize austerity policies and cost-cutting measures that maintain resource scarcity. Their introduction may be precisely because of policy changes and aligned reform movements, making it difficult to isolate the specific effects of the predictive tools themselves.For example, for risk assessment instruments (RAIs) in criminal justice, it is difficult to disentangle potential impacts of RAIs on reducing racial disparities from simultaneous braoder prosecutorial or bail reform movements. Conversely, attempts to enact reforms by computational changes to ADS are brittle: if front-line decision-makers disagree with the goals, they might simply ignore ADS information.
Example 1 (Risk assessment for pretrial decision-making).
When someone is charged with a criminal offense and placed in jail, judges decide whether the arrested person should be released before trial and under what conditions. Pretrial risk assessments aim to help judges make these decisions by predicting whether a defendant will fail to appear for future court dates or commit a crime if released. Although laws vary across U.S. jurisdictions, judges are typically required to weigh an individual’s risk of flight and future criminal activity when choosing from several pretrial options: unconditional release, release with non-monetary conditions like supervision or drug testing, money bail (which requires defendants to pay a set amount to secure release or remain in jail), or detention without the possibility of release. These consequential decisions often must be made quickly and with limited information about the defendant.
Pretrial risk assessments have been advanced as potential correctives for several long-standing issues in this challenging decision-making environment. By purporting to help judges more accurately assess a defendant’s risk, risk assessments have been suggested as tools to address over-detention and over-reliance on cash bail — widely criticized for tying release to financial means — as well as racial biases in decision-making and public safety concerns. A substantial body of scholarship and public writing has critiqued risk assessments on legal, ethical, and practical grounds (for reviews, see Eckhouse et al. (2018); Robinson and Koepke (2019)), yet they remain key components of reform efforts. Most notably, when New Jersey essentially eliminated cash bail statewide in 2017, it required judges to consider an algorithmic risk assessment in their decision-making while retaining judicial discretion over final pretrial determinations.
@@
Therefore, current decisions made in these organizations reflect a combination of institutional policy, categorical prioritization, and human discretion. In Figure 1, this is represented via the database of prior decisions and outcomes ($D$,$Y$).
@@
Institutional context and their ongoing practices introduce idiosyncratic variability in the requirements and goals of shared predictive tasks like “predict heart disease” or “predict recidivism”. A general task like “predict heart disease” becomes “predict heart disease when doctors and hospitals have protocol A,B,C for tests and medications, and the patient case mix is D.” We highlight case studies that illustrate how ADS differs from other prediction tasks.
@@
Institutional context dictates what data is collected in the first place—usually primarily for administrative purposes, rather than downstream research. [^4] For example, in Example 2 (cost prediction), a machine learning researcher might consider cost prediction rather than outcome prediction, as in (Obermeyer et al., 2019), to be an error of naivete rather than a natural consequence of how institutions generate data. A lot of data collection in healthcare is to support insurance billing, and was not collected in the first place to support downstream predictive innovation or future machine-learning papers.
@@ 3  Model Design
Although predictive risk models have a simple formulation—predict $Y$ from XX—they in fact are used in a wider variety of ways to impact and change future real-world outcomes.
ADS directs additional resources, scrutiny, services, protection/prevention, or sanction. These decisions often use individual-level information to make targeted decisions and need not be data-driven. These distinct use cases of ADS often imply different appropriate technical formulations beyond pure prediction. Later we discuss such prediction-intervention mismatches that arise in criminal justice (supervised release Example 4) and housing (Example 3).
@@
[…]
@@
While we use $D$ to represent historical decisions, we consider alternative decision-making procedures, summarized in the decision policy $\pi(X):\mathcal{X}\mapsto\mathcal{D}$. Predictive risk modeling often triages scarce resources. When used as such, often the decisions are thresholded based on predictive risk,
@@
In practice, we find that purely-predictive formulations that target on $\mathbb{E}[Y\mid X]$ are more common, i.e. learning a predictive model based on all prior data without any explicit modeling of prior decisions and interventions. However, if prior data does includes other decisions or interventions, this estimand hides the resulting heterogeneity, and is vulnerable to distribution shifts in interventions (including those across jurisdictions when an ADS developed in one is used in another). For example, a prior pilot study may not continue or might in fact scale up. Learning $\mathbb{E}[Y\mid X]$ from the pooled historical data would be brittle to such shifts, while the counterfactual risk models discussed above would remain robust [^7] Koepke and Robinson (2018) argue that ignoring recent reforms or other changes in the social world could lead to “zombie predictions” that overestimate risk. These interventions in the historical data include pretrial services that reduce rearrest risk or bail reforms that reduce pretrial detention and iatrogenic rearrests. .
@@
Intervention-aware targeting offers many different options beyond predictive risk models of $\mathbb{E}[Y\mid X]$ alone. Recent discussions of predictions vs. interventions include Barabas et al. (2018), which argues that the use of predictive risk models is a poor fit for improving social conditions, drawing on the case of risk assessments in criminal justice. [^8] Barabas et al. (2018) argues that predictive models that focus on merely predicting future outcomes neither manage long-term goals of reducing reoffense (especially when interactions with the criminal justice system are iatrogenic, increasing criminal activity), nor reduce criminal risk, nor improve understanding of root causes of criminal behavior that could be used to design interventions to reduce risk. This critique, among others, points out the urgency of going beyond prediction to consider alternative avenues for improving social outcomes. Prediction by itself doesn’t directly achieve long-term goals or inform interventions to reduce risk, as argued in Liu et al. (2023a). One can also reconceptualize predictive risk models in terms of the subjects of prediction. [^9] Meyer et al. (2022) highlight how current risk prediction focuses narrowly on harms individuals may pose to systems (e.g., future crime); by “flipping the script,” they show how prediction could just as well be used to predict system harms to individuals, such as lengthy sentences. So much more is possible beyond current formulations.
@@
Understanding the institutional context is key to inform prioritization for supervised release. Pretrial detention can preclude re-offense by its nature, albeit at a large cost to civil liberties. On the other hand, supervised release is a much “softer” decision and its effects on re-offense are unclear. How should judges determine who should receive supervised release? [^12] There is some broad guidance that “release conditions—if any are considered necessary—should be the least restrictive that reasonably assure court appearance and community safety,” (APPR, 2024) emphasizing $Y(D)$, the final outcomes of whatever release conditions are provided. Other guidance does indicate some preference for reducing risk of the riskiest individuals, suggesting that “[limited community] resources should go to people with the highest need, meaning people assessed as the least likely to succeed.” Heterogeneous treatment effects can introduce tradeoffs between these two competing objectives [^13] Increased surveillance can increase frequency of technical violations of the supervision conditions and therefore worsen outcomes. Supportive services might work for some on the margin, but may not be enough to overcome other structural barriers. The limited evidence base is mixed, and variation in its program design makes generalization from prior studies even more difficult. .
@@
$\text{recommended supervision}=\mathbb{I}\left[\mathbb{E}[Y^{FTA}(0)+Y^{NCA}(0)\mid X]>t\right],$
@@
We discuss details of the study in more detail in Appendix B. Even though the risk prediction model of Black et al. (2003) improved predictions compared to prior models in other states, in part by using additional covariates and using a continuous rather than binarized outcomes, they find that predicted duration is in general unrelated to heterogeneous treatment effects. In this setting, different equity vs. efficiency objectives might favor different problem formulations.
Other recent studies assess targeting based on baseline risk or need vs. welfare-efficient ITRs in different domains (Haushofer et al., 2022; Brown et al., 2016; Athey et al., 2025; Sirota et al., 2025).
Finally, we again acknowledge that there are also other important factors that justify targeting on $Y(0)$ rather than Y(1)Y(0)Y(1)-Y(0), such as strong moral emphasis on targeting the worst off, alignment with intuitive or politically palatable conceptions [^15] Gowan (2010) studies homelessness and policy developments in San Francisco; the medicalization of homelessness received greater institutional support and influenced the development of vulnerability assessments that focus on mortality risk. In turn, these vulnerability assessments individualize homelessness and emphasize individual-level health factors rather than broader structural and systemic factors contributing to homelessness., or simply that it is simpler and statistically easier to estimate $\mathbb{E}[Y(0)\mid X]$ rather than heterogeneous treatment effects $\mathbb{E}[Y(1)-Y(0)\mid X].$
@@
Formally, the first case is where we are able to specify a loss function $\ell(D,Y)$ which depends on an observable quantity $Y$ which is not itself a function of $D$, e.g., the decision to carry an umbrella does not itself impact whether it rains. The second case is where the appropriate loss function is parameterized by quantities that are not fully observable due to decision-dependence. For patients that we do not treat, we cannot observe their treatment effects, and vice versa.
[…]
@@
What are examples where prediction is enough? Liu et al. (2023a) formalized a notion of a sufficient action that improves outcomes for all decision subjects, regardless of their particular circumstances. Using a causal-graphical formalism, they show that the existence of such an action is both necessary and sufficient for pure prediction to be adequate for targeting interventions.
@@
Different strategies for learning decision policies include “plug-in prediction”, plugging in prediction models $\mathbb{E}[Y\mid X]$ into expected loss, optimization over decision rules, or more advanced “decision-aware learning” approaches that seek to learn predictive models that explicitly achieve the best decision-loss. An example of non-trivial decision-loss structure arises when cost minimization relies on global system state, such as shortest-paths, routing, and/or supply chain problems. See Appendix B for more details.
@@
In contrast, decisions impact outcomes in unknown ways (fig. 5(b)), targeting decisions requires estimating HTEs, as we have discussed before. For example, if our goal is to optimize post-decision outcomes $\min_{\pi\in\Pi}\mathbb{E}[Y(\pi(X))]$, then the implicit loss function $\ell(D,(Y(0),Y(1)))=Y(D)$ depends on both $Y(1)$ and $Y(0)$, which are never simultaneously observable.
@@
Although we have laid out basic distinctions between predictions and decisions, specifying the objective function and the constraints of the decision problem are also crucial. We refer the interested reader to (Barocas et al., 2023) and references therein for discussions of additional desiderata, such as algorithmic fairness. [^16] The algorithmic fairness literature has developed tools for formalizing certain desiderata (such as group or individual fairness), and achieving them, often with constrained optimization. In practice, while auditing for unfairness has surfaced harms and issues with deployed predictive models, sources of unfairness and appropriate mitigations vary widely: perhaps an inappropriate target variable was chosen, the data ecosystem introduces statistical bias, the prediction only informs an institutional decision that is missing from the problem formulation, and so on. A central critique of considerations in the algorithmic fairness literature—that is closely related to the theme in this work—is the mismatch between constraints on model predictions and actual utility or improvements from decisions (see e.g., Liu et al., 2018). Other concerns include “leveling down” objections to imposing fairness constraints that might worsen outcomes for one group while improving them for others (i.e. Pareto-suboptimal transfers of welfare) (Temkin, 2000; Mittelstadt et al., 2023; Hu and Chen, 2020; Maheshwari et al., 2023). Other perspectives on algorithmic fairness emphasize consequentialist or welfarist accounts (Corbett-Davies et al., 2017; Corbett-Davies and Goel, 2018; Hu and Chen, 2020), which are well aligned with our focus on evaluating realized impacts. Accurate and contextual measurement of realized impacts is a crucial input into any of these perspectives.
@@
While we have outlined different technical formulations, how does one translate a real-world context into a technical problem formulation? We discuss important organizational considerations in problem formulation. Next, we discuss what kind of domain knowledge is required. Finally, even if prediction is sufficient, our prior discussions suggest opportunities to leverage interventional analysis to improve the design of predictive risk models.
@@
In data-rich settings, it can be tempting to instead take a ‘data-driven’ opportunistic approach: exploring opportunities to innovate by training candidate predictive risk models on available outcomes in the data, and proceeding further if a model achieves a high enough accuracy. Such an approach, however, is liable to produce algorithms that are ill-suited for the contexts in which they will be used (Coston et al., 2023). Problem formulation should begin before looking at any data, involving key stakeholders in a conversation about the goals, success criteria, and potential risks of an algorithm in the proposed context (Kawakami et al., 2024). Efforts to improve the model design and development process would benefit from an understanding of how data science teams currently navigate the model development process (Passi and Barocas, 2019; Passi and Jackson, 2018; Guerdan et al., 2025; Kross and Guo, 2021).
The act of formalizing a problem and building a predictive model is in itself a policy intervention (Abebe et al., 2020). Formalizing a problem can enforce clarity around goals, but introducing a quantitative frame for prediction can distort multiple goals by elevating and operationalizing some but not others (Green and Chen, 2021; Johnson and Zhang, 2022; Liu et al., 2023b; Elzayn et al., 2024).
A key lever is the choice of which decisions or interventions to target. Causal analysis can identify key causal determinants of outcomes, and inform the choice of interventional frame. For example, Lee et al. (2023) studied the causal determinants of length of hospital stay after cardiac surgery. Clinicians were interested in which factors — such as medical staffing in the operating room — could be changed to reduce length of stay.